# Generating Structured Outputs from Language Models: Benchmark and Studies

### v1

@@ 1 Introduction
The rapid advancements in LMs in recent years have significantly broadened their applications, extending beyond natural language tasks to more complex challenges such as web navigation (Yao et al., 2023b), data extraction (Polak & Morgan, 2024), and tool use (Schick et al., 2023). Unlike traditional natural language processing (NLP) tasks where the output is aimed at review by humans, output in these applications is often consumed by machines such as controller and service APIs. The machine-oriented nature of these applications requires LMs to generate structured outputs that strictly adhere to predefined formats and constraints. However, the LM generation process is probabilistic and does not provide guarantees on the output’s structure, making it challenging to deploy LMs in applications requiring structured inputs and high reliability.
@@
The strong demand for structured generation (Liu et al., 2024) has led to the development of various constrained-decoding frameworks [^1] We use the terms constrained decoding framework and grammar engine interchangeably., such as Guidance (Guidance AI, 2023), Outlines (Willard & Louf, 2023), XGrammar (Dong et al., 2024) and the grammar module of Llamacpp (Gerganov & al., 2023) These frameworks provide broad support for different types of constraints, minimal overhead, and compatibility with various LM ecosystems, facilitating the adoption of constrained decoding in real-world applications.
@@
To answer these questions, we need to study constrained-decoding methods with a large-scale, diverse, and real-world collection of user-defined structures. To evaluate the performance of constrained decoding frameworks, we introduce JSONSchemaBench, a collection of 10K real-world JSON schemas from various sources, Organized into 10 datasets of varying complexity and diversity, the benchmark spans domains such as function signatures, service APIs, and system configurations. We evaluate six state-of-the-art constrained decoding frameworks, including Guidance, Outlines, Llamacpp, XGrammar, OpenAI, and Gemini, on JSONSchemaBench. We pair this real-world schema dataset with the official JSON Schema Test Suite (JSON Schema Org, 2024) in order to extract detailed insights into coverage of JSON Schema functionality across these frameworks, and to further evaluate them with considerations of end-to-end task accuracy in the context of multiple real-world tasks. Altogether, our evaluation takes three aspects into consideration: efficiency, coverage, and quality. We define specific metrics to measure these three functional aspects and evaluate constrained decoding frameworks against them. Through extensive experiments, we converge on the following findings as illustrated in Figure 1. (1) Constrained decoding can speed up the generation process by 50% compared to unconstrained decoding. (2) Frameworks demonstrate significant differences in their actual support for real-world JSON schemas, with the best framework supporting twice as many schemas as the worst. (3) Constrained decoding consistently improves the performance of downstream tasks up to 4%, even for tasks with minimal structure like GSM8k.
@@ 2 Background and Related Work
JSON Schema is a meta-language that describes the structure of JSON data. It is capable of expressing a wide variety of constraints, such as the types of JSON object properties, the length of JSON arrays or the pattern that a JSON string must match. The syntax and capabilities of JSON Schema are defined in the JSON Schema specification (Wright et al., 2022), which defines a large number of keywords, each of which may be used or combined with other keywords within a schema to enforce constraints like the ones mentioned. JSON Schema is widely used in the software ecosystem, and previous work has been done to collect extensive examples of JSON Schemas with a focus both on real-world use as well as on overall correctness.
Baazizi et al. (2021) collected over 6,000 JSON schemas from publicly available GitHub repositories. Attouche et al. (2022) used it alongside additional collected JSON schemas in order to evaluate a witness generation algorithm for JSON Schema. Separately, the official JSON Schema Test Suite (JSON Schema Org, 2024) is a collection of manually created test cases, maintained by the JSON Schema core team, which exercises a large portion of the functionality defined in the JSON Schema specification. It was originally written to assist implementers of JSON Schema validation tools with testing their compliance against the specification, and therefore contains a wide variety of examples for each of JSON Schema’s keywords, including in edge case scenarios. Notably, Bowtie (Bowtie, 2025) leverages the test suite as a foundation for comparing and understanding different implementations of the JSON Schema specification across programming languages. Taken together, these two datasets form a large number of examples both of JSON Schema’s diverse feature set as well as its use in the wild.
@@
Constrained decoding (Deutsch et al., 2019; Shin et al., 2021; Scholak et al., 2021; Poesia et al., 2022; Wang et al., ; Geng et al., 2023) refers to methods that guide the generation process of language models (LMs) by masking out tokens that do not adhere to predefined constraints at each step. Recently, highly optimized grammar-constrained decoding frameworks (Guidance AI, 2023; Beurer-Kellner et al., 2023; Willard & Louf, 2023; Kuchnik et al., 2023; Zheng et al., 2024; Dong et al., 2024) have been developed to improve the efficiency and usability of constrained decoding.
The evaluation of constrained decoding remains an under-explored topic, with no consensus on what defines the effectiveness of constrained decoding. While some research has pursued comparisons of constrained decoding with unconstrained LMs (Roy et al., 2024; Tang et al., 2024; Yao et al., 2023a), the studies to date fail to provide comparisons across different constrained decoding frameworks. The benchmarks employed have either narrowly focused on specific tasks or rely on formal-grammar–based artificial setups, that have unclear relevance to real-world use cases.
@@ 3.1 Data Collection
We start with the 6K JSON schemas collected by Baazizi et al. (2021) from publicly available GitHub repositories, and with the set of schemas from the JSON Schema Test Suite (JSON Schema Org, 2024). We further collect JSON schemas from other sources, such as the JSON Schema Store (Schema Store Org, 2014), the GlaiveAI function calling dataset V2 (GlaiveAI, 2024), and from Kubernetes configuration files (Kubernetes, 2022). We filter out invalid schemas and standardize the schemas to ensure that they conform to the version of JSON Schema declared [^2] The \$schema keyword, defined in the JSON Schema specification, allows any schema to self-identify which version of JSON Schema it is written for. in each schema
@@
Table 1: Schema collection metadata.
The GitHub JSON schemas collection from Baazizi et al. (2021) contains schemas of varying complexity and diversity, ranging from simple type constraints to complex constraints with nested objects and arrays. For more fine-grained evaluation, we split the data into five collections based on the schema size: trivial, small, medium, large, ultra. The suites finalized after all collection and processing are listed in Table 1. We excluded GitHub-Trivial and GitHub-Ultra from the experiments as they were too easy or too hard. However, we retained these datasets in the benchmark, with GitHub-Ultra serving as an aspirational target for future advancements. For more information on post-processing and dataset splitting, we refer the reader to Appendix A.
@@ 4 Efficiency
Naïve implementations of constrained decoding add overhead to the standard LM inference process, including a per-step mask computation and an optional one-time grammar compilation. However, several optimizations can significantly reduce this overhead. For instance, mask computation can run in parallel with the LM’s forward pass, and grammar compilation can be performed concurrently with pre-filling computations (Guidance AI, 2023; Dong et al., 2024). Other optimizations such as grammar caching and constraint-based speculative decoding (GuidanceAI, 2024b; Beurer-Kellner et al., 2023; Kurt, 2024a) can further reduce overhead.
@@ 4.1 Setup
While Outlines and Llamacpp demonstrate substantially lower throughput than the LM-only approach, Guidance achieves even higher efficiency, which it accomplishes by skipping certain generation steps with its guidance acceleration  (GuidanceAI, 2024b). Comparing Guidance and XGrammar with the HF Transformers backend shows that Guidance has a significantly better TPOT.
@@ 5.1 Setup
To measure empirical coverage, we conduct all experiments using the Llama-3.2-1B-Instruct model as it is small enough to run efficiently while still producing high-quality outputs. The prompt consists of a simple instruction with two-shot examples (Figure 3), and validation is performed using the jsonschema Python library (Berman (2025)) (using JSON Schema Draft2020-12) with string-format checks enabled. We use greedy decoding with zero-temperature, performing a single generation run, and enforce a 40-second timeout for grammar compilation and an additional 40 seconds for generation. Exceeding these limits is treated as a schema processing failure. Additional details are provided in Appendix B.
@@ 5.2 Results
Guidance shows the highest empirical coverage on six out of the eight datasets, with Llamacpp taking the lead on the remaining two: the domain-specific Washington Post and notably hard JSON Schema Store. On the other hand, closed-source grammar engines consistently have the lowest coverage; they came in last on all but one dataset. LM-only [^4] The Llama 3.1 models have been specifically fine-tuned to adhere to JSON schemas (Grattafiori et al., 2024) approaches achieve acceptable coverage on easy-to-medium datasets but show significant performance drops on harder datasets, such as Github Hard and JSON Schema Store, as well as domain-specific datasets like Washington Post. We note that while empirical coverage is a reasonable indicator of a framework’s real-world performance, it is influenced by factors such as the LM being used and the sampling methods employed.
@@ 5.3 JSON Schema Test Suite: Complementary Evaluation
Originally designed to test the correctness and compliance of JSON Schema validation implementations, the official JSON Schema Test Suite (JSON Schema Org, 2024) is a comprehensive collection of test cases spanning the many features of the JSON Schema specification. We believe that the test suite is an ideal tool for assessing the correctness of grammar engines.
@@ 6 Quality
In principle, constrained decoding should not affect the quality of the generated output as it only filters out the invalid tokens. However, things become more complicated due to ambiguity of tokenization (Vivien, 2024; GuidanceAI, 2024a; Geng et al., 2024) and the distributional shifts caused by the intervention (Geng et al., 2023; Tam et al., 2024). As a hypothetical toy example, an LM might answer ‘89,000’ instead of the correct ‘89000’ in a GSM8K question. Constrained decoding can block the invalid token ‘,’, enforcing structural compliance but potentially may cause the LM to go out of distribution and generate ‘890000’ instead. Kurt (2024b) argued that the performance decline observed in previous studies (Tam et al., 2024) comes from inadequate prompting, insufficient contextual information, and poorly crafted schemas.
@@ 6.1 Setup
Kurt (2024b); Tam et al. (2024) have introduced a series of tasks to investigate potential quality concerns in constrained decoding, which we leverage and extend in this benchmark. Specifically, we adopt the three reasoning tasks from these studies to evaluate the impact of constrained decoding on task accuracy, as detailed in Table LABEL:tab:task_description. The simple output structure of these tasks was designed to isolate the effects of constrained decoding on reasoning, as outlined by Tam et al. (2024).
For our experiments, we use the Llama-3.1-8B-Instruct model to measure task performance. We follow the original setup and prompt specifications from Kurt (2024b), with full details provided in Appendix F.
@@ 6.2 Results
The results in Table 8 show that the constrained decoding, regardless of the framework, achieves higher performance than the unconstrained setting. Among the frameworks evaluated, Guidance consistently delivers the best performance across all tasks, with approximately a 3% improvement over the LM-only approach in every task. We believe this may be attributed to its token-healing implementation (GuidanceAI, 2024a).
@@ Appendix A JSON Schema Collections Details
JSONSchemaBench includes a diverse collection of schemas curated from multiple real-world applicationsAttouche et al. (2022), designed to represent a wide range of use cases:
@@
  • GitHub (Baazizi et al., 2021): Extracted from open-source repositories containing schema definitions, representing practical, widely-used applications. Schemas from GitHub are of various complexities, totaling 6,000 schemas. We split the collection into trivial (fewer than 10 fields), easy (10–30 fields), medium (30–100 fields), hard (100–500 fields), and ultra (more than 500 fields), based on the total number of fields in each JSON schema to reflect increasing complexity and scale.
  • Snowplow (Analytics, 2022): Sourced from event-based analytics frameworks, showcasing schemas tailored for event-driven data structures.
  • Kubernetes (Kubernetes, 2022): Schemas defining configurations for container orchestration systems, highlighting schemas with intricate hierarchical structures.
  • WashingtonPost (Post, 2022): Schemas for The Washington Post’s ANS specification.
  • GlaiveAI2K GlaiveAI (2024): 2,000 schemas extracted from a function-calling dataset. Each schema represents a function signature.
  • JSON Schema Store (Schema Store Org, 2014): The largest collection of independent JSON schemas in the world.
@@
Table 9: Baisc statistics of the datasets used in the experiments.
@@ A.2 Draft versions
Table 10: JSON Schema Draft Version Counts
@@ Appendix E Efficiency Experiment Details
Dataset Framework GCT (s) TTFT (s) TPOT (ms) Washington Post LM only NA 0.14 16.54 Guidance 0.00 0.33 8.58 Llamacpp 0.06 0.24 27.20 Outlines 3.06 3.30 40.12 GitHub Hard LM only NA 0.51 16.18 Guidance 0.02 1.18 8.36 Llamacpp 0.05 0.60 28.17 Outlines 12.77 13.30 81.22
Table 13: This table provides additional results for Table 2 in the main body.

### v2

@@ 1 Introduction
The rapid advancements in LMs in recent years have significantly broadened their applications, extending beyond natural language tasks to more complex challenges such as web navigation [Yao et al., 2023b], data extraction [Polak & Morgan, 2024], and tool use [Schick et al., 2023]. Unlike traditional natural language processing (NLP) tasks where the output is aimed at review by humans, output in these applications is often consumed by machines such as controller and service APIs. The machine-oriented nature of these applications requires LMs to generate structured outputs that strictly adhere to predefined formats and constraints. However, the LM generation process is probabilistic and does not provide guarantees on the output’s structure, making it challenging to deploy LMs in applications requiring structured inputs and high reliability.
@@
The strong demand for structured generation [Liu et al., 2024] has led to the development of various constrained-decoding frameworks [^1] We use the terms constrained decoding framework and grammar engine interchangeably., such as Guidance [Guidance AI, 2023], Outlines [Willard & Louf, 2023], XGrammar [Dong et al., 2024] and the grammar module of Llamacpp [Gerganov & al., 2023] These frameworks provide broad support for different types of constraints, minimal overhead, and compatibility with various LM ecosystems, facilitating the adoption of constrained decoding in real-world applications.
@@
To answer these questions, we need to study constrained-decoding methods with a large-scale, diverse, and real-world collection of user-defined structures. To evaluate the performance of constrained decoding frameworks, we introduce JSONSchemaBench, a collection of 10K real-world JSON schemas from various sources, Organized into 10 datasets of varying complexity and diversity, the benchmark spans domains such as function signatures, service APIs, and system configurations. We evaluate six state-of-the-art constrained decoding frameworks, including Guidance, Outlines, Llamacpp, XGrammar, OpenAI, and Gemini, on JSONSchemaBench. We pair this real-world schema dataset with the official JSON Schema Test Suite [JSON Schema Org, 2024] in order to extract detailed insights into coverage of JSON Schema functionality across these frameworks, and to further evaluate them with considerations of end-to-end task accuracy in the context of multiple real-world tasks. Altogether, our evaluation takes three aspects into consideration: efficiency, coverage, and quality. We define specific metrics to measure these three functional aspects and evaluate constrained decoding frameworks against them. Through extensive experiments, we converge on the following findings as illustrated in Figure 1. (1) Constrained decoding can speed up the generation process by 50% compared to unconstrained decoding. (2) Frameworks demonstrate significant differences in their actual support for real-world JSON schemas, with the best framework supporting twice as many schemas as the worst. (3) Constrained decoding consistently improves the performance of downstream tasks up to 4%, even for tasks with minimal structure like GSM8k.
@@ 2 Background and Related Work
JSON Schema is a meta-language that describes the structure of JSON data. It is capable of expressing a wide variety of constraints, such as the types of JSON object properties, the length of JSON arrays or the pattern that a JSON string must match. The syntax and capabilities of JSON Schema are defined in the JSON Schema specification [Wright et al., 2022], which defines a large number of keywords, each of which may be used or combined with other keywords within a schema to enforce constraints like the ones mentioned. JSON Schema is widely used in the software ecosystem, and previous work has been done to collect extensive examples of JSON Schemas with a focus both on real-world use as well as on overall correctness.
Baazizi et al. [2021] collected over 6,000 JSON schemas from publicly available GitHub repositories. Attouche et al. [2022] used it alongside additional collected JSON schemas in order to evaluate a witness generation algorithm for JSON Schema. Separately, the official JSON Schema Test Suite [JSON Schema Org, 2024] is a collection of manually created test cases, maintained by the JSON Schema core team, which exercises a large portion of the functionality defined in the JSON Schema specification. It was originally written to assist implementers of JSON Schema validation tools with testing their compliance against the specification, and therefore contains a wide variety of examples for each of JSON Schema’s keywords, including in edge case scenarios. Notably, Bowtie [Bowtie, 2025] leverages the test suite as a foundation for comparing and understanding different implementations of the JSON Schema specification across programming languages. Taken together, these two datasets form a large number of examples both of JSON Schema’s diverse feature set as well as its use in the wild.
@@
Constrained decoding [Deutsch et al., 2019; Shin et al., 2021; Scholak et al., 2021; Poesia et al., 2022; Wang et al., ; Geng et al., 2023] refers to methods that guide the generation process of language models (LMs) by masking out tokens that do not adhere to predefined constraints at each step. Recently, highly optimized grammar-constrained decoding frameworks [Guidance AI, 2023; Beurer-Kellner et al., 2023; Willard & Louf, 2023; Kuchnik et al., 2023; Zheng et al., 2024; Dong et al., 2024] have been developed to improve the efficiency and usability of constrained decoding.
The evaluation of constrained decoding remains an under-explored topic, with no consensus on what defines the effectiveness of constrained decoding. While some research has pursued comparisons of constrained decoding with unconstrained LMs [Roy et al., 2024; Tang et al., 2024; Yao et al., 2023a], the studies to date fail to provide comparisons across different constrained decoding frameworks. The benchmarks employed have either narrowly focused on specific tasks or rely on formal-grammar–based artificial setups, that have unclear relevance to real-world use cases.
@@ 3.1 Data Collection
We start with the 6K JSON schemas collected by Baazizi et al. [2021] from publicly available GitHub repositories, and with the set of schemas from the JSON Schema Test Suite [JSON Schema Org, 2024]. We further collect JSON schemas from other sources, such as the JSON Schema Store [Schema Store Org, 2014], the GlaiveAI function calling dataset V2 [GlaiveAI, 2024], and from Kubernetes configuration files [Kubernetes, 2022]. We filter out invalid schemas and standardize the schemas to ensure that they conform to the version of JSON Schema declared [^2] The \$schema keyword, defined in the JSON Schema specification, allows any schema to self-identify which version of JSON Schema it is written for. in each schema
Table 1: Schema collection metadata.
@@
The GitHub JSON schemas collection from Baazizi et al. [2021] contains schemas of varying complexity and diversity, ranging from simple type constraints to complex constraints with nested objects and arrays. For more fine-grained evaluation, we split the data into five collections based on the schema size: trivial, small, medium, large, ultra. The suites finalized after all collection and processing are listed in Table 1. We excluded GitHub-Trivial and GitHub-Ultra from the experiments as they were too easy or too hard. However, we retained these datasets in the benchmark, with GitHub-Ultra serving as an aspirational target for future advancements. For more information on post-processing and dataset splitting, we refer the reader to Appendix A.
@@ 4 Efficiency
Naïve implementations of constrained decoding add overhead to the standard LM inference process, including a per-step mask computation and an optional one-time grammar compilation. However, several optimizations can significantly reduce this overhead. For instance, mask computation can run in parallel with the LM’s forward pass, and grammar compilation can be performed concurrently with pre-filling computations [Guidance AI, 2023; Dong et al., 2024]. Other optimizations such as grammar caching and constraint-based speculative decoding [GuidanceAI, 2024b; Beurer-Kellner et al., 2023; Kurt, 2024a] can further reduce overhead.
@@ 4.1 Setup
While Outlines and Llamacpp demonstrate substantially lower throughput than the LM-only approach, Guidance achieves even higher efficiency, which it accomplishes by fast-forwarding [^3] See Tables13 and 14 for the number of tokens fast-forwarded in the experiments. certain generation steps with its guidance acceleration  [GuidanceAI, 2024b]. Comparing Guidance and XGrammar with the HF Transformers backend shows that Guidance has a significantly better TPOT.
@@ 5.1 Setup
To measure empirical coverage, we conduct all experiments using the Llama-3.2-1B-Instruct model as it is small enough to run efficiently while still producing high-quality outputs. The prompt consists of a simple instruction with two-shot examples (Figure 3), and validation is performed using the jsonschema Python library (Berman [2025]) (using JSON Schema Draft2020-12) with string-format checks enabled. We use greedy decoding with zero-temperature, performing a single generation run, and enforce a 40-second timeout for grammar compilation and an additional 40 seconds for generation. Exceeding these limits is treated as a schema processing failure. Additional details are provided in Appendix B.
@@ 5.2 Results
Guidance shows the highest empirical coverage on six out of the eight datasets, with Llamacpp taking the lead on the remaining two: the domain-specific Washington Post and notably hard JSON Schema Store. On the other hand, closed-source grammar engines consistently have the lowest coverage; they came in last on all but one dataset. LM-only [^5] The Llama 3.1 models have been specifically fine-tuned to adhere to JSON schemas [Grattafiori et al., 2024] approaches achieve acceptable coverage on easy-to-medium datasets but show significant performance drops on harder datasets, such as Github Hard and JSON Schema Store, as well as domain-specific datasets like Washington Post. We note that while empirical coverage is a reasonable indicator of a framework’s real-world performance, it is influenced by factors such as the LM being used and the sampling methods employed.
@@ 5.3 JSON Schema Test Suite: Complementary Evaluation
Originally designed to test the correctness and compliance of JSON Schema validation implementations, the official JSON Schema Test Suite [JSON Schema Org, 2024] is a comprehensive collection of test cases spanning the many features of the JSON Schema specification. We believe that the test suite is an ideal tool for assessing the correctness of grammar engines.
@@ 6 Quality
In principle, constrained decoding should not affect the quality of the generated output as it only filters out the invalid tokens. However, things become more complicated due to ambiguity of tokenization [Vivien, 2024; GuidanceAI, 2024a; Geng et al., 2024] and the distributional shifts caused by the intervention [Geng et al., 2023; Tam et al., 2024]. As a hypothetical toy example, an LM might answer ‘89,000’ instead of the correct ‘89000’ in a GSM8K question. Constrained decoding can block the invalid token ‘,’, enforcing structural compliance but potentially may cause the LM to go out of distribution and generate ‘890000’ instead. Kurt [2024b] argued that the performance decline observed in previous studies [Tam et al., 2024] comes from inadequate prompting, insufficient contextual information, and poorly crafted schemas.
@@ 6.1 Setup
Kurt [2024b]; Tam et al. [2024] have introduced a series of tasks to investigate potential quality concerns in constrained decoding, which we leverage and extend in this benchmark. Specifically, we adopt the three reasoning tasks from these studies to evaluate the impact of constrained decoding on task accuracy, as detailed in Table LABEL:tab:task_description. The simple output structure of these tasks was designed to isolate the effects of constrained decoding on reasoning, as outlined by Tam et al. [2024].
For our experiments, we use the Llama-3.1-8B-Instruct model to measure task performance. We follow the original setup and prompt specifications from Kurt [2024b], with full details provided in Appendix F.
@@ 6.2 Results
The results in Table 8 show that the constrained decoding, regardless of the framework, achieves higher performance than the unconstrained setting. Among the frameworks evaluated, Guidance consistently delivers the best performance across all tasks, with approximately a 3% improvement over the LM-only approach in every task. We believe this may be attributed to its token-healing implementation [GuidanceAI, 2024a].
@@ Appendix A JSON Schema Collections Details
JSONSchemaBench includes a diverse collection of schemas curated from multiple real-world applicationsAttouche et al. [2022], designed to represent a wide range of use cases:
@@
  • GitHub [Baazizi et al., 2021]: Extracted from open-source repositories containing schema definitions, representing practical, widely-used applications. Schemas from GitHub are of various complexities, totaling 6,000 schemas. We split the collection into trivial (fewer than 10 fields), easy (10–30 fields), medium (30–100 fields), hard (100–500 fields), and ultra (more than 500 fields), based on the total number of fields in each JSON schema to reflect increasing complexity and scale.
  • Snowplow [Analytics, 2022]: Sourced from event-based analytics frameworks, showcasing schemas tailored for event-driven data structures.
  • Kubernetes [Kubernetes, 2022]: Schemas defining configurations for container orchestration systems, highlighting schemas with intricate hierarchical structures.
  • WashingtonPost [Post, 2022]: Schemas for The Washington Post’s ANS specification.
  • GlaiveAI2K GlaiveAI [2024]: 2,000 schemas extracted from a function-calling dataset. Each schema represents a function signature.
  • JSON Schema Store [Schema Store Org, 2014]: The largest collection of independent JSON schemas in the world.
Table 9: Baisc statistics of the datasets used in the experiments.
@@ A.2 Draft versions
Table 10: JSON Schema Draft Version Counts
@@ Appendix E Efficiency Experiment Details
Table 13: Efficiency metrics for different engines with LlamaCpp as the inference engine. GCT: Grammar Compilation Time, TTFT: Time to First Token, TPOT: Time Per Output Token, TGT: Total Generation Time, FF: Fast-Forwarded output tokens. Bold values indicate the smallest in each column for GCT, TTFT, TPOT, and TGT. All values are median of the samples.
Dataset Framework GCT (s) TTFT (s) TPOT (ms) TGT (s) Output Tokens (FF) GlaiveAI LLM only NA 0.10 15.40 1.08 64.94 (00.00) Guidance 0.00 0.24 6.37 0.50 41.56 (15.70) Llamacpp 0.05 0.20 29.98 1.47 43.18 (00.00) Outlines 3.48 3.65 30.33 4.84 40.39 (00.00) GitHub Easy LLM only NA 0.10 15.83 0.95 53.91 (00.00) Guidance 0.00 0.34 7.44 0.60 34.92 (10.02) Llamacpp 0.05 0.18 27.22 1.10 33.93 (00.00) Outlines 3.71 3.97 39.78 5.29 34.19 (00.00) Snowplow LLM only NA 0.11 16.23 1.01 55.31 (00.00) Guidance 0.00 0.28 6.55 0.51 36.77 (14.50) Llamacpp 0.05 0.20 28.90 1.24 37.21 (00.00) Outlines 3.91 4.14 42.66 5.65 35.65 (00.00) GitHub Medium LLM only NA 0.20 16.68 2.56 142.10 (00.00) Guidance 0.01 0.54 7.57 1.29 99.66 (31.42) Llamacpp 0.06 0.30 29.08 2.85 87.71 (00.00) Outlines 8.05 8.38 46.57 12.23 84.64 (00.00) Kubernetes LLM only NA 0.16 15.32 0.84 44.38 (00.00) Guidance 0.01 0.45 9.47 0.71 28.75 (04.40) Llamacpp 0.05 0.28 28.04 1.06 28.09 (00.00) Outlines 5.29 5.55 46.10 6.56 22.26 (00.00)
Table 14: Efficiency metrics for different engines with Hugging Face Transformers as the inference engine. All values are median of the samples.
Dataset Framework GCT (s) TTFT (s) TPOT (ms) TGT (s) Output Tokens (FF) GlaiveAI Guidance 0.01 0.36 36.92 1.87 41.45(16.76) XGrammar 0.12 0.30 66.78 2.87 39.47(00.00) GitHub Easy Guidance 0.01 0.37 42.03 1.60 27.67(06.75) XGrammar 0.11 0.33 65.57 4.07 59.45(00.00) GitHub Medium Guidance 0.01 0.55 44.21 4.84 96.31(26.93) XGrammar 0.20 0.48 65.51 6.53 92.93(00.00) GitHub Hard Guidance 0.01 0.73 35.88 10.25 211.40(101.40) XGrammar 0.30 0.65 65.20 14.99 221.40(00.00)